### v3

@@
Abstract. The literature on cluster-randomized trials typically assumes no interference across clusters. This may be implausible when units are irregularly distributed in space without well-separated communities, in which case clusters may not represent significant geographic, social, or economic divisions. In this paper, we develop methods for reducing bias due to cross-cluster interference. First, we propose an estimation strategy that excludes units not surrounded by clusters assigned to the same treatment arm. We show that this substantially reduces asymptotic bias relative to conventional difference-in-means estimators without substantial cost to variance. Second, we formally establish a bias-variance trade-off in the choice of clusters: constructing fewer, larger clusters reduces bias due to interference but increases variance. We provide a rule for choosing the number of clusters to balance the asymptotic orders of the bias and variance of our estimator. Finally, we consider unsupervised learning for cluster construction and provide theoretical guarantees for $k$-medoids.
@@ 1 Introduction
CRTs conducted on large populations have become increasingly common in academia (Muralidharan and Niehaus, 2017) and industry (Karrer et al., 2021). For example, Egger et al. (2022) study the general equilibrium effects of unconditional cash transfers through a CRT conducted in rural Kenya. In their trial, villages are grouped into contiguous clusters called “sublocations,” the basic administrative unit above the village. The authors observe that “villages are relatively close to each other [and] sublocation boundaries are not ‘hard’ in any sense nor reflective of salient ethnic or social divides…there is extensive economic interaction in nearby markets regardless of sublocation.”
@@
This paper proposes methods for reducing bias due to cross-cluster interference through the analysis and design of cluster-randomized trials. Instead of partial interference, we consider a spatial interference model proposed by Leung (2022b) which posits that interference decays with geographic distance. This captures the central concern expressed in the previous examples, potential interference between geographically proximate units.
@@
At the analysis stage, a common bias-reduction strategy is the “fried-egg design” (Hayes and Moulton, 2017, Ch. 4.4.3). This is not an experimental design but rather entails excluding from estimation units in the trial located near cluster boundaries, the “whites” of the “fried eggs” that are the clusters. An open question is whether excluding observations in such a fashion is worth the loss of efficiency. We show that conventional difference-in-means estimators can have large biases due to interference near cluster boundaries. We then propose to improve the efficiency of the fried-egg design by excluding not all units near cluster boundaries but rather only the subset not surrounded by clusters assigned to the same treatment arm. We prove that this can substantially reduce the asymptotic order of the bias relative to difference in means with no increase to the asymptotic order of the variance.
We then turn to optimal design of clusters. For motivation, consider the typhoid vaccine trial studied by Sur et al. (2009). They conduct a CRT in an urban slum in India spanning only about 1.2 by 0.7 km, dividing 38k participants into 80 contiguous clusters. By comparison, the previously referenced Homan et al. (2016) trial has about the same number of participants and clusters, but theirs takes place in a substantially larger region spanning about 12 by 4 km. While malaria and typhoid are distinct diseases, a concern is that the potential for cross-cluster interference is greater when unit density is higher.
@@
The fundamental statistical problem is a bias-variance trade-off in the number of clusters. Partitioning the region into fewer, larger clusters reduces bias from cross-cluster interference since fewer units are located near cluster boundaries, the regions most prone to bias due to interference. However, this comes at the cost of higher variance because the sample size in a CRT is the number of clusters. Unlike the conventional partial interference framework, our spatial interference model generates the bias-variance trade-off, which enables the study of cluster design for optimizing the trade-off.
@@
We prove that the number of clusters $k$ that balances the asymptotic bias and variance of our estimators depends on a parameter $\gamma$ that measures the speed at which interference decays with distance. This formally characterizes how domain knowledge of interference informs optimal cluster construction, knowledge that is implicitly used in practice when researchers define fried-egg boundaries or construct clusters to minimize interaction between units.
@@
In practice, several ad hoc methods are used to construct clusters. Homan et al. (2016) utilize a “traveling salesman algorithm,” essentially an unsupervised learning method. More commonly, researchers use administrative divisions (e.g. Egger et al., 2022) or manually partition the study region (e.g. Moulton et al., 2001). Binka et al. (1998) study a malaria trial conducted in Northern Ghana, noting that, “Where possible, small paths or road were used to delineate the clusters. However, in most cases, the cluster boundaries did not correspond to natural barriers.” We contribute to the literature by providing formal justification for constructing clusters using $k$-medoids, a well-known unsupervised learning algorithm.
@@
Our work is most closely related to Leung (2022b). He studies designs targeting the global average treatment effect (GATE) in which clusters are squares in $\mathbb{R}^{2}$ with identical areas. We target a more general set of estimands and study richer designs that may utilize unsupervised learning to construct clusters. We also consider more general metric spaces allowing for “infill-increasing” settings in which the number of units is of larger asymptotic order than the volume of the study region. This accommodates relatively “denser” settings such as that of Sur et al. (2009). Finally, we propose new standard errors that are asymptotically conservative without restrictions on the superpopulation.
Our analysis of $k$-medoids builds on Cao et al. (2022). We extend a key result of theirs to the case where $k$ diverges and characterize other $k$-medoid properties in this regime which may be of independent interest. We generalize their Ahlfours-regularity condition to allow for the infill-increasing case, though unlike them, we also require an additional boundary condition.
Faridani and Niehaus (2023) substantially generalize the theoretical results in Leung (2022b) while retaining his focus on the GATE. Their results hold for general Euclidean-like spaces, which they define using topological conditions that differ from Ahlfours regularity. For this reason, their proofs differ substantially from ours. They study designs that essentially cover the study region with non-intersecting balls of approximate radius $g_{n}$, where the radius is chosen to grow at an optimal rate. Our design specifies an optimal choice for the number of clusters and uses unsupervised learning construct clusters.
The paper is organized as follows. The next section defines the model and estimands. In § 3, we discuss the disadvantages with existing estimators and propose our alternative. We study their theoretical properties in § 4 and propose an optimal design in § 5. We present simulation results in § 6 and an empirical application in § 7 using data from the Egger et al. (2022) trial. Finally, § 8 concludes and summarizes the practical outputs of our analysis.
@@ 2 Setup
Let $\mathcal{X}$ be a set of spatial locations and $\rho$ a metric on $\mathcal{X}$. We observe a set of $n$ units $\mathcal{N}_{n}\subseteq\mathcal{X}$, so $\rho(i,j)$ is the spatial distance between units $i,j\in\mathcal{N}_{n}$. Let $D_{i}$ denote unit $i$’s binary treatment assignment and $\bm{D}=(D_{i})_{i\in\mathcal{N}_{n}}$ the vector of observed assignments, which is the only random quantity in our analysis. Let $\{Y_{i}(\cdot)\}_{i\in\mathcal{N}_{n}}$ be a set of functions with domain $\{0,1\}^{n}$ and range $\mathbb{R}$. For $\bm{d}=(d_{i})_{i\in\mathcal{N}_{n}}\in\{0,1\}^{n}$, $Y_{i}(\bm{d})$ denotes the potential outcome of unit $i$ under the counterfactual that all units are assigned treatments according to $\bm{d}$. Unit $i$’s observed outcome is $Y_{i}=Y_{i}(\bm{D})$. We maintain the following standard assumption, that potential outcomes are uniformly asymptotically bounded (all asymptotic statements are with respect to $n\in\mathbb{N}$ unless otherwise indicated).
@@ 2.1 Spatial Interference
We next formally specify the metric space in which units are located. The leading example of interest is $\mathbb{R}^{2}$, but our results apply to a more general “Ahlfors-regular” space, augmented with an additional boundary condition. Define the $r$-neighborhood of unit $i$
@@
The spatial dimension is $d$, while $\xi_{n}$ is a measure of density (units per unit volume). The dependence on $\xi_{n}$ is new to this paper relative to the prior work on spatial interference and accommodates applications with denser regions. To understand the assumption, consider the standard increasing-domain case in which $\xi_{n}=1$ for all $n$. Part (a) says that the number of units in an $r$-neighborhood is the same order $r^{d}$ as the neighborhood’s volume. This defines a $(C,d)$-finite Ahlfors-regular space, the same metric space studied by Cao et al. (2022). The boundary condition in part (b) is new relative to their setup, but both (b) and the upper bound in (a) are satisfied if $\mathcal{X}=\mathbb{R}^{d}$ under the usual increasing-domain assumption that units are minimally separated in space (Jenish and Prucha, 2009, Lemma A.1).
Assumption 2 also accommodates the infill-increasing case in which $\xi_{n}$ diverges with $n$. Here the number of units in any neighborhood is of larger order than the neighborhood volume by a factor of $\xi_{n}$, corresponding to a denser region. Because we require $\xi_{n}\prec n$, this is a hybrid of infill and increasing-domain asymptotics in which the volume of the study region $\mathcal{N}_{n}$ still grows but at a slower rate $n/\xi_{n}$ than the population size $n$ (as in Lahiri and Zhu, 2006, for example).
Under Assumption 2(a), the number of units in any $r$-ball is proportional to $\xi_{n}r^{d}$, which allows for some variation in density across the study region. However, if two subregions of similar volume exist but the number of units in one is a multiple of the other that is large relative to the sample size, then the assumption is violated. This could be the case in practice if the study region encompasses urban and rural areas. In § SA.2.1, we discuss how our proposed methodology can be modified to accommodate large variations in density.
@@
This corresponds to Assumption 3 of Leung (2022b). Unlike partial interference, it is formulated independently of clusters, enabling us to develop a theory of optimal cluster construction. To interpret the condition, consider a unit $i$ centered at a “cluster” of radius $r$. The quantity $\lvert Y_{i}(\bm{d})-Y_{i}(\bm{d}^{\prime})\rvert$ measures interference induced by manipulating the treatment assignments of units outside the cluster, and the assumption requires this to decay like $r^{-\gamma}$ or faster. Hence, the larger the minimum distance $r$ between $i$ and the units with manipulated treatments, the smaller the spillover effect. We thus interpret $1/\gamma$ as (an upper bound on) the degree of interference.
Assumption 3 further requires $\gamma$ to decay fast enough relative to the spatial dimension $d$. This coincides with the requirement imposed by Leung (2022b) for $d=2$. In the increasing-domain case, the intuition is that $r$-neighborhood sizes grow like $r^{d}$ under Assumption 2, so weak dependence requires interference to decay faster than the rate at which neighborhoods densify with $r$. Central limit theorems for spatial processes impose analogous requirements on mixing coefficients, which control the degree of spatial dependence as $\gamma$ does in our framework (e.g. Jenish and Prucha, 2009, Assumption 3(b)).
@@ 2.2 Design
Part (a) says that clusters are globular in that they contain and are contained by balls with radii of the same order. Under Assumption 2(a), part (b) implies that the number of units in any cluster is order $n/k_{n}$, which implies that clusters are comparable in size. We show below that clusters generated by $k$-medoids satisfy these requirements. Note that even under partial interference, restrictions on cluster size heterogeneity are required for inference using conventional clustered standard errors (Hansen and Lee, 2019, Assumption 2), although our requirements are stronger. While these may be possible to relax, we leave this to future research.
Figure 1 depicts clusters theoretically guaranteed to satisfy Assumption 5, illustrating that the conditions allow clusters to have quite heterogeneous shapes. However, the assumption is violated if elongated clusters exist that narrowly encompass a geographical feature such as a river or road or if clusters are highly size-imbalanced. These issues may be addressed by grouping adjacent clusters or subdividing large or abnormally shaped clusters. The assumption can also be violated if clusters are not constructed based on spatial proximity. For instance, if units are connected through an online social network in which units far apart in space are well connected, clusters based on network connectivity would likely violate the assumption.
Assumption 5 is satisfied if the researcher partitions the space into $k_{n}$ identically-sized cubes, but such clusters do not adapt to the spatial distribution of units. A better alternative is to use unsupervised learning algorithms to partition space. We next provide theoretical guarantees for the well-known $k$-medoids algorithm, which is stated in § SA.1.
@@
Suppose clusters are the output of the $k$-medoids procedure given in Algorithm SA.1.1. Under Assumption 2(a), the clusters satisfy Assumption 5.
In practice, $k$-means can also be used since it typically delivers clusters similar to those of $k$-medoids. The globular clusters produced by $k$-medoids and $k$-means are sometimes viewed unfavorably relative to algorithms such as spectral clustering for certain unsupervised learning tasks. However, for our purposes, globular clusters are preferable because they minimize the number of units near cluster boundaries, which are the primary source of bias due to cross-cluster interference.
@@ 2.3 Causal Estimands
Because we allow for cross-cluster interference, we need to redefine conventional estimands in a manner free of this source of bias. To this end, we define an idealized counterfactual design distinct from the “factual” design in Assumption 4. For $t\in\{0,1\}$, the $p_{t}$-single-cluster design sets $k=1$ and $q=t$ in Assumption 4, so that the entire population is grouped into a single cluster and assigned to treatment with probability $p_{t}$. Throughout the paper, let ${\bf E}[\cdot]$ denote the expectation taken with respect to the factual design and ${\bf E}^{*}_{p_{t}}[\cdot]$ the expectation taken with respect to the $p_{t}$-single-cluster design.
@@
for $d_{1},d_{0}\in\{\varnothing,0,1\}$, where we define $Y_{i}(\varnothing,\bm{D}_{-i})\equiv Y_{i}(\bm{D})$. The following special cases are analogous to estimands defined by Hudgens and Halloran (2008) and Hayes and Moulton (2017). The direct effect compares treated and untreated units under the $p_{1}$-single-cluster design:
@@
This is directly identified if the $p_{1}$-single-cluster design were implemented in practice. The estimands that follow, however, require multiple clusters assigned to different arms because they compare different counterfactual designs. Under our framework, this is the primary motivation for cluster randomization.
@@
Interest often centers on the “pure control” baseline of $p_{0}=0$, so that ${\bf E}^{*}_{p_{0}}[Y_{i}(0,\bm{D}_{-i})]=Y_{i}(\bm{0})$. The total effect is the sum of the direct and indirect effects, equal to $\theta_{T}^{*}=\theta^{*}(1,0;p_{1},p_{0})$. Finally, the overall effect compares outcomes under different saturation levels: $\theta_{O}^{*}=\theta^{*}(\varnothing,\varnothing,p_{1},p_{0})$.

### v4

@@
Abstract. The literature on cluster-randomized trials typically allows for interference within but not across clusters. This may be implausible when units are irregularly distributed across space without well-separated communities, as clusters in such cases may not align with significant geographic, social, or economic divisions. This paper develops methods for reducing bias due to cross-cluster interference. We first propose an estimation strategy that excludes units not surrounded by clusters assigned to the same treatment arm. We show that this substantially reduces bias relative to conventional difference-in-means estimators without significant cost to variance. Second, we formally establish a bias-variance trade-off in the choice of clusters: constructing fewer, larger clusters reduces bias due to interference but increases variance. We provide a rule for choosing the number of clusters to balance the asymptotic orders of the bias and variance of our estimator. Finally, we consider unsupervised learning for cluster construction and provide theoretical guarantees for $k$-medoids.
@@ 1 Introduction
CRTs conducted on large populations have become increasingly common in academia (Muralidharan and Niehaus, 2017) and industry (Karrer et al., 2021). For example, Egger et al. (2022) study the general equilibrium effects of unconditional cash transfers using a CRT conducted in rural Kenya. The trial clustered villages into administrative units called “sublocations.” The authors observe that “villages are relatively close to each other [and] sublocation boundaries are not ‘hard’ in any sense nor reflective of salient ethnic or social divides…there is extensive economic interaction in nearby markets regardless of sublocation.”
@@
This paper proposes methods for reducing bias due to cross-cluster interference in the analysis and design of cluster-randomized trials. Instead of partial interference, we consider a spatial interference model proposed by Leung (2022b) which posits that interference decays with geographic distance. This captures the central concern expressed in the previous examples, namely potential interference between geographically proximate units.
@@
At the analysis stage, a common bias-reduction strategy is the “fried-egg design” (Hayes and Moulton, 2017, Ch. 4.4.3). This is not an experimental design but rather entails excluding from estimation units in the trial located near cluster boundaries, the “whites” of the “fried eggs” that are the clusters. An open question is whether excluding observations in such a fashion is worth the loss of efficiency. We show that conventional difference-in-means estimators can have large biases due to interference near cluster boundaries. We then propose to improve the efficiency of the fried-egg design by excluding not all units near cluster boundaries but rather only the subset not surrounded by clusters assigned to the same treatment arm. We prove that this can substantially reduce the asymptotic order of the bias relative to difference in means with no increase in the asymptotic order of the variance.
@@
We then turn to optimal design of clusters. Partitioning the region into a small number of large clusters reduces bias since fewer units are located near cluster boundaries where they are most prone to cross-cluster interference. However, this comes at the cost of higher variance because the sample size in a CRT is the number of clusters. Unlike the conventional partial interference framework, our spatial interference model induces a bias-variance trade-off, which enables the study of cluster design for optimizing the trade-off.
@@
We prove that the number of clusters $k$ that balances the asymptotic bias and variance of our estimators depends on a parameter $\gamma$ that measures the speed at which interference decays with distance. This formally characterizes how domain knowledge of interference informs optimal cluster construction. Such knowledge is implicitly used in practice when researchers define fried-egg boundaries or construct clusters with “buffer zones” to minimize interaction between units (Hayes and Moulton, 2017).
@@
In practice, cluster construction is ad hoc. Homan et al. (2016) utilize a “traveling salesman algorithm,” essentially an unsupervised learning method. More commonly, researchers use administrative divisions (e.g. Egger et al., 2022) or manually partition the study region (e.g. Moulton et al., 2001). Binka et al. (1998) study a malaria trial conducted in Northern Ghana, noting that, “Where possible, small paths or road were used to delineate the clusters. However, in most cases, the cluster boundaries did not correspond to natural barriers.” We contribute to the literature by providing formal justification for constructing clusters using $k$-medoids, a well-known unsupervised learning algorithm.
@@
Our work is most closely related to Leung (2022b). He studies designs targeting the global average treatment effect (GATE) in which clusters are squares in $\mathbb{R}^{2}$ with identical areas. We consider a more general set of estimands and richer designs that may utilize unsupervised learning to construct clusters. Also, to accommodate denser spatial settings, we consider “infill-increasing” asymptotics in which the number of units is of larger asymptotic order than the volume of the study region. Finally, we propose new standard errors that are asymptotically conservative without restrictions on the superpopulation.
Our analysis of $k$-medoids builds on Cao et al. (2024). We extend a key result of theirs to the case where $k$ diverges and characterize other $k$-medoid properties in this regime which may be of independent interest. We generalize their Ahlfours-regularity condition on the metric space to allow for infill-increasing asymptotics, though unlike them, we also require an additional boundary condition.
Faridani and Niehaus (2024) substantially generalize the theoretical results in Leung (2022b) while retaining his focus on the GATE. Their results hold for general spaces defined using topological conditions that differ from Ahlfours regularity. For this reason, their proofs differ substantially from ours. They study designs that essentially cover the study region with non-intersecting balls of approximate radius $g_{n}$, where the radius is chosen to grow at an optimal rate. Our design specifies an optimal choice for the number of clusters and uses unsupervised learning to construct clusters.
The paper is organized as follows. The next section defines the model and estimands. In § 3, we discuss the disadvantages of existing estimators and propose an alternative. We study the theoretical properties of the estimators in § 4 and propose an optimal design in § 5. We present simulation results in § 6 and an empirical application in § 7 using data from the Egger et al. (2022) trial. Finally, § 8 concludes and summarizes the practical outputs of our analysis.
@@ 2 Setup
Let $(\mathcal{X},\rho)$ be a metric space where $\mathcal{X}$ is the set of spatial locations and $\rho$ the metric. We observe a set of $n$ units $\mathcal{N}_{n}\subseteq\mathcal{X}$, so $\rho(i,j)$ is the spatial distance between units $i,j\in\mathcal{N}_{n}$. Let $D_{i}$ denote unit $i$’s binary treatment assignment and $\bm{D}=(D_{i})_{i\in\mathcal{N}_{n}}$ the vector of observed assignments, which is the only random quantity in our analysis. Let $\{Y_{i}(\cdot)\}_{i\in\mathcal{N}_{n}}$ be a set of functions with domain $\{0,1\}^{n}$ and range $\mathbb{R}$. For $\bm{d}=(d_{i})_{i\in\mathcal{N}_{n}}\in\{0,1\}^{n}$, $Y_{i}(\bm{d})$ denotes the potential outcome of unit $i$ under the counterfactual that all units are assigned treatments according to $\bm{d}$. Unit $i$’s observed outcome is $Y_{i}=Y_{i}(\bm{D})$. We maintain the following standard assumption, that potential outcomes are uniformly asymptotically bounded. All asymptotic statements are with respect to a sequence indexed by $n\in\mathbb{N}$ unless otherwise indicated.
@@ 2.1 Spatial Interference
The primary metric space of interest is $\mathbb{R}^{2}$, but our results apply to a more general “Ahlfors-regular” space, augmented with a boundary condition. Define the $r$-neighborhood of unit $i$
@@
The constant $d$ represents the spatial dimension and $\xi_{n}$ a measure of density (units per unit volume). The dependence on $\xi_{n}$ is new to this paper relative to prior work on spatial interference and accommodates applications with denser regions. To understand the assumption, consider the standard increasing-domain case in which $\xi_{n}=1$ for all $n$. Part (a) says that the number of units in an $r$-neighborhood is the same order $r^{d}$ as the neighborhood’s volume. This defines a $(C,d)$-finite Ahlfors-regular space, the same metric space studied by Cao et al. (2024). The boundary condition in part (b) is new relative to their setup, but both (b) and the upper bound in (a) are satisfied if $\mathcal{X}=\mathbb{R}^{d}$ under the usual increasing-domain assumption that units are minimally separated in space (Jenish and Prucha, 2009, Lemma A.1).
Assumption 2 also accommodates the infill-increasing case in which $\xi_{n}$ diverges with $n$. Here the number of units in any neighborhood is of larger order than the neighborhood volume by a factor of $\xi_{n}$, corresponding to a denser region. Because we require $\xi_{n}\prec n$, this is a hybrid of infill and increasing-domain asymptotics in which the volume of the study region $\mathcal{N}_{n}$ grows but at a slower rate $n/\xi_{n}$ than the population size $n$ (as in Lahiri and Zhu, 2006, for example).
Under Assumption 2(a), the number of units in any $r$-ball is proportional to $\xi_{n}r^{d}$, which allows for some variation in density across the study region. The assumption is violated if two subregions of similar volume exist but the number of units in one is a large multiple of the other. This could be the case in practice if the study region encompasses urban and rural areas. In § SA.2.1, we discuss how our proposed methodology can be modified to accommodate large variations in density.
@@
This corresponds to Assumption 3 of Leung (2022b). Unlike partial interference, it is formulated independently of clusters, enabling us to develop a theory of optimal cluster construction. To interpret the condition, consider a unit $i$ centered at a “cluster” $\mathcal{N}(i,r)$ of radius $r$. The quantity $\lvert Y_{i}(\bm{d})-Y_{i}(\bm{d}^{\prime})\rvert$ measures interference induced by manipulating the treatment assignments of units outside the cluster, and the assumption requires this to decay like $r^{-\gamma}$ or faster. Hence, the larger the minimum distance $r$ between $i$ and the units with manipulated treatments, the smaller the spillover effect. We thus interpret $1/\gamma$ as (an upper bound on) the degree of interference. In § 5, we discuss the optimal design of clusters using knowledge of $\gamma$.
Assumption 3 further requires $\gamma$ to decay fast enough relative to the spatial dimension $d$. This coincides with the requirement imposed by Leung (2022b) for $d=2$. In the increasing-domain case, the intuition is that $r$-neighborhood sizes grow like $r^{d}$ under Assumption 2, so weak dependence requires interference to decay faster than the rate at which neighborhoods densify with $r$. Central limit theorems for spatial processes impose analogous requirements on mixing coefficients which control the degree of spatial dependence, as $\gamma$ does in our framework (e.g. Jenish and Prucha, 2009, Assumption 3(b)).
@@ 2.2 Design
This requires clusters to be globular in that they contain and are contained by balls with radii of the same order. Under Assumption 2(a), part (b) implies that the number of units in any cluster is order $n/k_{n}$, which implies that clusters are comparable in size. We show below that clusters generated by $k$-medoids satisfy these requirements. Note that even under partial interference, restrictions on cluster size heterogeneity are required for inference using conventional clustered standard errors (Hansen and Lee, 2019, Assumption 2), although our requirements are stronger. These may be possible to relax, but we leave this to future research.
Assumption 5 allows clusters to have quite heterogeneous shapes, as can be seen in Figure 1. It depicts $k$-medoid clusters, which satisfy Assumption 5 by Theorem 1 below. However, the assumption is violated if clusters are highly size-imbalanced, such as elongated clusters that narrowly encompass a geographical feature such as a river or road. This may be addressed by subdividing large or abnormally shaped clusters or grouping adjacent small clusters. The assumption can also be violated if clusters are not constructed based on spatial proximity. For instance, if units are connected through an online social network connecting units far apart in space, then clusters based on network connectivity would likely violate the assumption.
Assumption 5 is satisfied if the researcher partitions the space into $k_{n}$ identically-sized cubes, but such clusters do not adapt to the spatial distribution of units. We therefore suggest using unsupervised learning algorithms. The next result provides theoretical guarantees for the well-known $k$-medoids algorithm stated in § SA.1.
@@
Suppose clusters are the output of $k$-medoids, given in Algorithm SA.1.1. Under Assumption 2(a), the clusters satisfy Assumption 5.
In practice, $k$-means can also be used since it typically delivers clusters similar to those of $k$-medoids. The globular clusters produced by both algorithms are sometimes viewed unfavorably relative to algorithms such as spectral clustering for certain unsupervised learning tasks. However, for our purposes, globular clusters are preferable because they minimize the number of units near cluster boundaries, which are the primary source of bias due to cross-cluster interference. More broadly, Assumption 5 provides general design principles for constructing clusters, namely to aim for balance and globularity, which may be achieved either using these algorithms or manually.
@@ 2.3 Causal Estimands
Because we allow for cross-cluster interference, we need to redefine conventional estimands in a manner free of this source of bias. To this end, define for $t\in\{0,1\}$ the $p_{t}$-counterfactual design, which sets $k=1$ and $q=t$ in Assumption 4. This design groups the population into a single cluster and assigns treatment with probability $p_{t}$. Throughout the paper, let ${\bf E}[\cdot]$ denote the expectation taken with respect to the observed design in Assumption 4 and ${\bf E}^{*}_{p_{t}}[\cdot]$ the expectation taken with respect to the $p_{t}$-counterfactual design.
@@
for $d_{1},d_{0}\in\{\varnothing,0,1\}$, where we define $Y_{i}(\varnothing,\bm{D}_{-i})\equiv Y_{i}(\bm{D})$. The following special cases are analogous to estimands defined by Hudgens and Halloran (2008) and Hayes and Moulton (2017). The direct effect compares treated and untreated units under the $p_{1}$-counterfactual design:
@@
This is directly identified if the $p_{1}$-counterfactual design were implemented in practice. The estimands that follow, however, require multiple clusters assigned to different arms because they compare different counterfactual designs. Under our framework, this is the primary motivation for cluster randomization.
@@
Interest often centers on the “pure control” baseline of $p_{0}=0$, so that ${\bf E}^{*}_{p_{0}}[Y_{i}(0,\bm{D}_{-i})]=Y_{i}(\bm{0})$. The total effect is the sum of the direct and indirect effects, equal to $\theta_{T}^{*}=\theta^{*}(1,0;p_{1},p_{0})$. Finally, the overall effect compares outcomes under different saturation levels: $\theta_{O}^{*}=\theta^{*}(\varnothing,\varnothing,p_{1},p_{0})$. Leung (2024) provides conditions under which these have causal interpretations.